| |
ID |
Date |
Author |
Subject |
Text |
 |
|
|
725
|
Fri Oct 24 14:33:37 2025 |
JB, GB, MP, AM | Snout grounding, bias and test input layout as of 24.10.2025 | |
| |
|
724
|
Fri Oct 24 12:29:02 2025 |
JB, GB, MP, AM, NH | 207Bi source threshold tests | > > >
> > > We have dismounted the AIDA stack and
bPlast and realigned the cables so that they |
11x | |
|
723
|
Fri Oct 24 10:31:59 2025 |
JB, GB, MP, AM, NH | 207Bi source threshold tests | > >
> > We have dismounted the AIDA stack and
bPlast and realigned the cables so that they |
6x | |
|
722
|
Thu Oct 23 18:45:12 2025 |
TD | Thursday 23 October 2025 |
19.45 Remote system check
|
6x | |
|
721
|
Thu Oct 23 12:22:31 2025 |
JB, GB, MP, AM, NH | 207Bi source threshold tests | >
> We have dismounted the AIDA stack and bPlast
and realigned the cables so that they can |
8x | |
|
720
|
Wed Oct 15 09:41:24 2025 |
JB, GB, MP | 207Bi source threshold tests |
We have dismounted the AIDA stack and bPlast
and realigned the cables so that they can |
  | |
|
719
|
Mon Oct 13 15:04:53 2025 |
JPB, NH, GB, MP | Dismount AIDA and realign the cables | We have dismounted the AIDA stack and bPlast
and realigned the cables so that they can
fit into the adaptors.
|
 | |
|
718
|
Thu Jul 24 15:50:01 2025 |
CC, MP, NK | Check of BB7 impact | HT of upstream AIDA from iseg, HT
cable shielded, 100V, 17 uA. HT
of downstream AIDA from CAEN, AIDA#0. Both |
   | |
|
717
|
Thu Jul 24 13:36:34 2025 |
CC, MP, NK | AIDA05 connected and disconnected compared to AIDA09 | HT cable shielded, HT
of upstream AIDA from iseg, HT of downstream
AIDA from CAEN. Both channel 0 and 1 of CAEN |
10x | |
|
716
|
Thu Jul 24 10:28:04 2025 |
CC, NK, MP | Comparison of channels in the same FEE64 and ASIC 05.1 and with AIDA13.1.8L | BB7 is still connected but not biased (same
configuration as the previous elog). The
channel of the CAEN power supply #0 is connected |
8x | |
|
715
|
Wed Jul 23 17:02:22 2025 |
CC, NK | Comparison of channels in the same FEE64 and ASIC 05.1 and with AIDA13.1.8L | BB7 is still connected but not biased (same
configuration as the previous elog). The
channel of the CAEN power supply #0 is connected |
8x | |
|
714
|
Wed Jul 23 15:56:00 2025 |
CC, MP, NK | Test of AIDA05 with new bias module | Attachment1: AIDA05.1.8L when AIDA05 connected,
biased at 101 V with shielded cables. BB7
is still connected but not biased (same configuration |
| |
|
713
|
Wed Jul 23 15:04:26 2025 |
CC, MP, NK | Test of AIDA05 with BB7 connected to Mesytec but not biased | Attachment1: AIDA05.1.8L when AIDA05 connected,
biased at 100 V with not-shielded cables,
after the DEGAS platform has been closed, |
| |
|
712
|
Wed Jul 23 13:59:37 2025 |
CC, NK | Shielding of the snout | Attachment1: AIDA05.1.8L when AIDA05 connected,
biased at 100 V with not-shielded cables,
CAEN HT, after the base of the |
| |
|
711
|
Wed Jul 23 13:08:30 2025 |
CC, NK | Comparison of channels in the same FEE64 and ASIC 05.1 | Biased at 100, AIDA05 connected, HT cables
not shielded CAEN HT
Attachment 1: AIDA05.1.8L --> |
    | |
|
710
|
Wed Jul 23 09:37:34 2025 |
CC, NK | Comparison pulser FWHM with/without HT and detector cables | This elog is the first reliable.
Before, the histograms have not been zero'd
Each test are done and compared |
9x | |
|
709
|
Tue Jul 22 18:00:45 2025 |
CC, NK, MP | FWHM aida FEE | AIDA05 completely disconnected, all the
others FEEs64 are fully connected
Attachment 1: aida05, FWHM= 95 |
 | |
|
708
|
Tue Jul 22 17:00:26 2025 |
MP, CC | Status | Test with shielded HV cables for both detectors.
The FWHM for ch1 is now 184 ch.
Yesterday cables for cards 5, 6, |
 | |
|
707
|
Mon Jul 21 13:01:19 2025 |
MP | Status | Status before grounding checks |
 | |
|
706
|
Sun Jun 15 10:53:08 2025 |
TD | Sunday 15 June 2025 |
11.45 DSSSD bias & leakage current - attachment
1
|
41x |
|